OpenStudy (anonymous):

An apartment building contains 12 units consisting of one- and two bedroom apartments that rent for $360 and $450 per month, respectively. When all units are rented the total monthly rental is $4950, what is the number of two bredroom apartments? A. 3 B.4 C. 5 D. 6 E. 7

OpenStudy (dogzcatz):

E

OpenStudy (anonymous):

how you know that @DogzCatz

OpenStudy (dogzcatz):

360x+5400-450x=4950 -90x=-450 x=5 y=12-5=7

OpenStudy (anonymous):

where did you get 5400 from?